OpenStudy (anonymous):

The equation is: (sqrt[2]/2-sqrt[2]/2i)+(-sqrt[2]/2-sqrt[2]/2i) Sorry if that's hard to understand. x_x The i represents an imaginary number. It's easier if you write the equation down on paper, probably.

OpenStudy (anonymous):

\[(\frac{\sqrt2}{2}-\frac{\sqrt2}{2}i)+(-\frac{\sqrt2}{2}-\frac{\sqrt2}{2}i)\]?

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

This is not an equation.

OpenStudy (anonymous):

Do you mean to simplify it?

OpenStudy (anonymous):

What would it be considered?

OpenStudy (anonymous):

Yes. Sorry about the confusion.

OpenStudy (anonymous):

\[(\frac{\sqrt2}{2}-\frac{\sqrt2}{2}i)+(-\frac{\sqrt2}{2}-\frac{\sqrt2}{2}i)\]\[=\frac{\sqrt2}{2}-\frac{\sqrt2}{2}i-\frac{\sqrt2}{2}-\frac{\sqrt2}{2}i\]\[=\frac{\sqrt2}{2}-\frac{\sqrt2}{2}-\frac{\sqrt2}{2}i-\frac{\sqrt2}{2}i\]So, do the subtraction for the first two terms first, can you?

OpenStudy (anonymous):

Would it be 0?

OpenStudy (anonymous):

Yes! So, you get \[=0-\frac{\sqrt2}{2}i-\frac{\sqrt2}{2}i\]Can you simplify the last two terms now?

OpenStudy (anonymous):

Wouldn't that be 0? Or am I doing something wrong?

OpenStudy (anonymous):

No What is -a -a?

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

No :S What is -1 -1?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Yup :) What is -2 - 2?

OpenStudy (anonymous):

-4!

OpenStudy (anonymous):

Yup! So what is -100 -100?

OpenStudy (anonymous):

-200.

OpenStudy (anonymous):

Yup, so what is -a -a ?

OpenStudy (anonymous):

-a?

OpenStudy (anonymous):

No :S -1 - 1 = - (1+1) = -2 -2 - 2 = - (2+2) = -4 -100 - 100 = - (100 + 100) = -200 -a - a = ...?

OpenStudy (anonymous):

-a - a -a + -a right?

OpenStudy (anonymous):

Yes. Which is?

OpenStudy (anonymous):

is a+a=2a?

OpenStudy (anonymous):

Yes What is -a - a?

OpenStudy (anonymous):

-2a?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Now, remember we've got\[=0-\frac{\sqrt2}{2}i-\frac{\sqrt2}{2}i\] We treated \(\frac{\sqrt2}{2}i\) = a, so it becomes = 0-a -a = ...?

OpenStudy (anonymous):

0-2a? Which would be, -2a. right?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Now, put back a = \(\frac{\sqrt2}{2}i\) into -2a, what do you get?

OpenStudy (anonymous):

I think I'm having trouble with the square roots. How do I simplify them?

OpenStudy (anonymous):

\[=-2 \times \frac{\sqrt2}{2}i = -\frac{2\sqrt2}{2}i\]There is something that you can cancel.

OpenStudy (anonymous):

The 2's?

OpenStudy (anonymous):

That aren't under the square root.

OpenStudy (anonymous):

Yes :) What do you get then?

OpenStudy (anonymous):

\[\sqrt{2}\]i

OpenStudy (anonymous):

It's supposed to be next to the 2.

OpenStudy (anonymous):

\[\sqrt2i\]Yes!

OpenStudy (anonymous):

but with a negative sign!

OpenStudy (anonymous):

\[=-2 \times \frac{\sqrt2}{2}i = -\frac{2\sqrt2}{2}i=-\frac{\cancel{2}\sqrt2}{\cancel{2}}i=-\sqrt2i\]